6.“Burgerina” is a country situated on the distant planet Jupiter. It produces only 2 types of burger and the  following information is presented to you:

In the year of 2016 the price & quantity of cheeseburgers -$4.00, 120

In the year of 2016 the price & quantity of Chickenburgers- $8.00, 10

In the year of 2017 the price & quantity of cheeseburgers- $5.00, 115

In the year of 2017 the price & quantity of Chickenburgers- $7.90, 7

In the year of 2018 the price & quantity of cheeseburgers- $8.00, 110

In the year of 2018 the price & quantity of Chickenburgers- $7.75, 2

a) Calculate nominal and real GDP (based on 2016) for Burgerina.

3. Based on the information presented in question 6, calculate the inflation rate for Burgerina for the years 2017 and 2018. Inflation should be calculated compared to previous years.


1
Expert's answer
2021-11-09T10:45:32-0500

Solution:

a.). Nominal GDP Year 2016 = (4 "\times" 120) + (8 "\times" 10) = 480 + 80 = 560

Nominal GDP Year 2017 = (5 "\times" 115) + (7.90 "\times" 7) = 575 + 55.3 = 630.3

Nominal GDP Year 2018 = (8 "\times" 110) + (7.75 "\times" 2) = 880 + 15.5 = 895.5

 

Real GDP Year 2016 = (4 "\times"120) + (8 "\times"10) = 480 + 80 = 560

Real GDP Year 2017 = (4 "\times" 115) + (8 "\times" 7) = 460 + 56 = 516

Real GDP Year 2018 = (4 "\times" 110) + (8 "\times" 2) = 440 + 16 = 456

 

3.). GDP Deflator = Nominal GDP/Real GDP

GDP Deflator 2016 = "\frac{560}{560} \times 100 = 100%"

GDP Deflator 2017 = "\frac{630.3}{516} \times 100 = 122\%"


GDP Deflator 2018 = "\frac{895.5}{456} \times 100 = 196\%" 895.5/456 x 100 = 196%

 

Inflation rate 2017 = (122% - 100%) = 22%

Inflation rate 2018 = (196% - 100%) = 96%